# A Peek into Token Bias: Large Language Models Are Not Yet Genuine Reasoners

### v1

@@ Abstract
This study introduces a hypothesis-testing framework to assess whether large language models (LLMs) possess genuine reasoning abilities or primarily depend on token bias. We go beyond evaluating LLMs on accuracy; rather, we aim to investigate their token bias in solving logical reasoning tasks. Specifically, we develop carefully controlled synthetic datasets, featuring conjunction fallacy and syllogistic problems. Our framework outlines a list of hypotheses where token biases are readily identifiable, with all null hypotheses assuming genuine reasoning capabilities of LLMs. The findings in this study suggest, with statistical guarantee, that most LLMs still struggle with logical reasoning. While they may perform well on classic problems, their success largely depends on recognizing superficial patterns with strong token bias, thereby raising concerns about their actual reasoning and generalization abilities. Codes are open-sourced at https://github.com/bowen-upenn/llm_token_bias.
@@ A Peek into Token Bias: Large Language Models Are Not Yet
Bowen Jiang1, Yangxinyu Xie1, 2, Zhuoqun Hao1, Xiaomeng Wang1, Tanwi Mallick2, Weijie J. Su1, Camillo J. Taylor1, Dan Roth1    University of Pennsylvania1   Argonne National Laboratory2 Philadelphia, PA, 19104, USA   Lemont, IL, 60439, USA {bwjiang@seas, xinyux@wharton, zhuoqunh@sas, xwang1@wharton}.upenn.edu, tmallick@anl.gov, {suw@wharton, cjtaylor@seas, danroth@seas}.upenn.edu
@@ 1 Introduction
Large language models (LLMs) have achieved remarkable progress in understanding and generating human-like text, triggering growing interest in the LLMs’ theory of minds Kosinski (2023); Jamali et al. (2023); Bubeck et al. (2023) and decision-making abilities Lyu et al. (2023); Prasad et al. (2023). However, there is ongoing debate about whether LLMs possess genuine reasoning capabilities, as evidence suggests that the performance of LLMs on reasoning tasks is correlated with how much the input’s semantic content supports a correct logical inference Dasgupta et al. (2022); Li et al. (2023). Should valid reasoning be applied, such a correlation would not exist, since a genuine reasoner should be able to derive the correct inference regardless of the context.
In this paper, we formalize this observation and say that an LLM is subject to token bias in a reasoning task if systematic changes to some or all tokens in the task descriptions - while keeping the underlying logic intact - allow us to predict the direction of the shift in the model’s output. A strong token bias suggests that the model is relying on superficial patterns in the input rather than truly understanding the underlying reasoning task. This could lead to brittle performance that fails to generalize well to novel examples and phrasings encountered in the wild that differ from the spurious patterns that the model has learned from the training data.
@@
Figure 2: What is token bias? Here is an example exhibited by GPT-4. On the left, GPT-4 correctly identifies the conjunction fallacy and answers the question correctly, given the classical Linda Problem as the one-shot exemplar. On the right, however, the exemplar is rephrased by altering "Linda" to "Bob" while keeping the same logic, which surprisingly confuses the model.
We explore several well-known logical fallacy problems from the cognitive science literature Tversky and Kahneman (1983); Kahneman (2011), which provide a clear playground for assessing the reasoning capabilities of LLMs. Figure 2 depicts one kind of token bias found in our testing framework, where the model may be overfitting to specific tokens commonly found in classic problem statements. Since we observe many cases where state-of-the-art LLMs like GPT-4 Achiam et al. (2023) successfully identify logical fallacies under certain settings, we highlight the urgent need for a framework to tease out whether LLMs apply genuine reasoning or merely exploit token bias for their improved performance.
@@
Our study is unique from existing work Gou et al. (2023); Suri et al. (2024); Mukherjee and Chang (2024); Wang et al. (2024) in two folds. First, we are not evaluating the overall accuracy of LLMs in identifying different logical fallacies. Instead, our focus is on token bias. Although there are always more types of logical fallacies, we take the conjunction fallacy and syllogistic fallacy as quintessential examples, which exhibit strong token biases that are more readily identifiable in their problem statements. By identifying and perturbing these specific tokens, we can induce predictable shifts in LLM responses. Second, we recognize that cognitive biases often emerge in implicit forms in real-life scenarios, so relying on engineering fine-grained prompts  Gou et al. (2023); Yao et al. (2024); Besta et al. (2024) to make LLMs identify specific logical fallacies is impractical for general-purpose user applications. As a result, we only leverage common prompting techniques that are sufficient to provide robust statistical evidence.
@@ 2 The General Framework
Once the underlying logic of a reasoning task is defined, we create an algorithm to generate a synthetic dataset with $n$ samples. While it is helpful to leverage LLMs for linguistic coherence in the process, the data generation should be carefully controlled, utilizing information from real-world data or established datasets to mitigate potential biases from purely AI-generated texts. The process begins with the creation of a curated list of entities and a textual template that dictates the structure of the task description. By sampling from this list, we generate task descriptions that maintain the integrity and novelty of the dataset. This method ensures that while the LLM of interest might be familiar with the individual entities, it has never seen the specific combinations of these entities and narratives, thus bypassing the risk of data contamination.
@@ 3.1 Preliminaries
The syllogistic fallacy documents the logical failure that occurs when people are presented with syllogisms – two premises followed by a conclusion. Ideally, if the premises are true and the logical structure is valid, the conclusion must necessarily be true. However, when the argument’s structure is flawed, the conclusion may be invalid despite the surface-level logical form. For example, consider the following syllogism from Kahneman (2011):
Is this logically sound?
@@ 3.3 Token bias on widely cited examples in classic literature
3.4 Token bias on celebrity names
@@ 3.6 Leaking Hint Tokens
We also manually craft a detailed chain-of-thought instructions Wei et al. (2022) that teach LLMs about correct reasoning steps and potential logical pitfalls, as shown in Appendix B.
@@ 4.2 Synthetic Dataset Generation
(c) Experimental results for Hypothesis 3.4 ($n=100$). The perturbed problems alternate the celebrity name to a generic one in problem statements. We run all different prompt methods. To reject the null, we expect ${\color[rgb]{0.921875,0.6796875,0.63671875}n12}<{\color[rgb]{0.28515625,0.5703125,0.6171875}n21}.$ We conclude that LLMs are frequently misled by irrelevant celebrity names in problem statements that are irrelevant to logical essence.
@@ 4.3 Prompting Methods
We implemented commonly used prompting strategies that are sufficient for evaluating the null hypotheses within our framework. The specific prompting techniques we utilized are as follows, with their corresponding notations presented in Figure 3: Baseline: Directly answering the question without additional instructions. Zero-shot chain-of-thought (zs_cot): Includes the instruction "Let us think step by step" Wei et al. (2022). One-shot (os): Involves a single in-context learning example Brown et al. (2020). Few-shots (fs): Utilizes three in-context examples. Similarly, we have os_cot and fs_cot. We also include weak_control_zs/os_cot and control_zs/os_cot for Hypothesis 3.6 representing prompts with additional weak or strong hints, as detailed in Appendix B.
@@ 5 Related Work
Recent studies Gardner et al. (2020); Hagendorff et al. (2023); Lin and Ng (2023); Talboy and Fuller (2023); Binz and Schulz (2023); Ullman (2023); Mitchell and Krakauer (2023) that propose synthetic datasets to analyze the biases in LLMs. For example, Tamkin et al. (2023) uses an LLM to generate potential prompts that reveals patterns of both positive and negative discrimination in LLMs. Echterhoff et al. (2024) proposes a set of LLM-simulated experiments in the context of college admissions to evaluate anchoring, framing, group attribution, and primacy bias. Although existing works Mukherjee and Chang (2024); Macmillan-Scott and Musolesi (2024); Wang et al. (2024); Suri et al. (2024) study more kinds of fallacy types in human psychology, they approach problems at a coarse level and only emphasize accuracy. Our study goes into a more fine-grained level with a series of hypotheses. We provide statistical guarantee and quantitative analysis of token bias that can be carefully tuned in a systematical way. Besides, Gou et al. (2023) presents the Rationality of Thought (RoT), decomposing responses into six predefined steps with hand-crafted prompt engineerings. Our work focuses on general prompting strategies that are sufficient to validate or reject our hypotheses.
@@ 6 Discussion
This work reconceptualizes the evaluation of the reasoning behavior of LLMs through the lens of token bias. The statistical evidence presented in our hypothesis-testing framework contributes to the larger discussion that LLMs do not apply reasoning consistently in their decision-making processes. Instead, they primarily rely on token bias for response generation. This suggests that chain-of-thought prompting Wei et al. (2022); Wang et al. (2022) or in-context learning Brown et al. (2020); Min et al. (2022); Lyu et al. (2022); Wang et al. (2022) may not elicit actual reasoning but instead result in semantic shortcuts for LLMs to imitate the desired behavior at superficial levels. These findings raise concerns about the extent to which LLMs truly engage in reasoning. Further investigations are needed to uncover the underlying mechanisms and limitations of LLMs’ reasoning capabilities.
@@ 7 Limitations
This hypothesis-testing framework is specifically designed for multiple-choice or yes/no questions and is not applicable to open-ended responses. It relies on LLMs with strong instruction-following capabilities to consistently produce responses that include the selected options, but we find that LLMs can generally follow these instructions in most cases. In addition, smaller LLMs, such as llama-3-8b-instruct, with lower instruction following capabilities may contain more confounders besides the token bias, which could weaken our hypothesis testing results. As a result, we mainly focus on state-of-the-art LLMs. Moreover, we acknowledge that there are likely other hypotheses and assumptions that a genuine reasoner should satisfy. Our current study focuses on the conjunction fallacy, syllogistic fallacy, and their variants to demonstrate our framework. These are quintessential examples and the framework could include a broader range of hypotheses, fallacy types, and reasoning tasks.
@@ C.1 Conjunction Fallacy
Suppose Joe Biden is running for president in 2024. Which is more likely:
(a) Joe Biden will win the national popular vote
(b) Joe Biden will win the national popular vote but lose the Electoral College vote
@@ D.1 Hypothesis 1
The full experimental results for Hypothesis 3.2 are shown in Figure 4 and Table LABEL:table:h1.
Table 2: Full Experimental results for Hypothesis 3.2

### v2

@@ Abstract
This study introduces a hypothesis-testing framework to assess whether large language models (LLMs) possess genuine reasoning abilities or primarily depend on token bias. We go beyond evaluating LLMs on accuracy; rather, we aim to investigate their token bias in solving logical reasoning tasks. Specifically, we develop carefully controlled synthetic datasets, featuring conjunction fallacy and syllogistic problems. Our framework outlines a list of hypotheses where token biases are readily identifiable, with all null hypotheses assuming genuine reasoning capabilities of LLMs. The findings in this study suggest, with statistical guarantee, that most LLMs still struggle with logical reasoning. While they may perform well on classic problems, their success largely depends on recognizing superficial patterns with strong token bias, thereby raising concerns about their actual reasoning and generalization abilities. Codes and data are open-sourced at https://github.com/bowen-upenn/llm_token_bias.
@@ A Peek into Token Bias: Large Language Models Are Not Yet
Bowen Jiang1, 2, Yangxinyu Xie1, 2, Zhuoqun Hao1, Xiaomeng Wang1, Tanwi Mallick2, Weijie J. Su1, Camillo J. Taylor1, Dan Roth1    University of Pennsylvania1   Argonne National Laboratory2 Philadelphia, PA, 19104, USA   Lemont, IL, 60439, USA {bwjiang@seas, xinyux@wharton, zhuoqunh@sas, xwang1@wharton}.upenn.edu, tmallick@anl.gov, {suw@wharton, cjtaylor@seas, danroth@seas}.upenn.edu
@@ 1 Introduction
Large language models (LLMs) have achieved remarkable progress in understanding and generating human-like text, triggering growing interest in the LLMs’ theory of minds Kosinski (2023); Jamali et al. (2023); Bubeck et al. (2023) and decision-making abilities Lyu et al. (2023); Prasad et al. (2023); Jiang et al. (2024a, b); Xie et al. (2024). However, there is ongoing debate about whether LLMs possess genuine reasoning capabilities, as evidence suggests that the performance of LLMs on reasoning tasks is correlated with how much the input’s semantic content supports a correct logical inference Dasgupta et al. (2022); Li et al. (2023). Should valid reasoning be applied, such a correlation would not exist, since a genuine reasoner should be able to derive the correct inference regardless of the context.
In this paper, we formalize this observation and say that an LLM is subject to token bias in a reasoning task if systematic changes to some or all tokens in the task descriptions - while keeping the underlying logic intact - allow us to predict the direction of the shift in the model’s output. A strong token bias suggests that the model is relying on superficial patterns in the input rather than truly understanding the underlying reasoning task. This could lead to brittle performance that fails to generalize well to novel examples and phrasings encountered in the wild, which could differ from the spurious patterns the model may have overfitted to during training.
Figure 1: We illustrate token bias using the classic "twenty-five horses" problem in graph theory. The top two sub-figures, generated by GPT-4o for illustration purposes only [^2] Interestingly, when we prompted GPT-4o to generate an image of "lop-eared bunnies", the model exhibited a visual token bias by depicting bunnies with four ears — both lop and erect — implying it associated the term "bunnies" with the presence of erect ears in images, without a genuine and logical understanding of a bunny’s physical reality. We move the exploration of visual token bias to the future work., demonstrate the concept by altering the name "horses" to "bunnies", irrelevant to the problem’s underlying logic. The bottom two sub-figures show experimental results in GPT-4 and Claude, where performance significantly drops due to perturbations in animal names and numbers. In these plots, "Original" refers to the unaltered "twenty-five horses" problem, "random_animals" alters only the animal names, and "random" alters both names and numbers. We observe ${\color[rgb]{0.921875,0.6796875,0.63671875}n12}>{\color[rgb]{0.28515625,0.5703125,0.6171875}n21}$ with statistical significance, meaning that there are more instances where the original problem is solved correctly while the perturbed problem is solved incorrectly, compared to the reverse. As a result, our hypothesis testing confirms token bias in this scenario.
@@
Figure 3: What is token bias? Here is another example exhibited by GPT-4. On the left, GPT-4 correctly identifies the conjunction fallacy and answers the question correctly, given the classical Linda Problem as the one-shot exemplar. On the right, however, the exemplar is rephrased by altering "Linda" to "Bob" while keeping the same logic, which surprisingly confuses the model.
We explore several well-known logical fallacy problems from the cognitive science literature Tversky and Kahneman (1983); Kahneman (2011), which provide a clear playground for assessing the reasoning capabilities of LLMs. Figure 2 and 3 depict two kinds of token biases found in our testing framework, where the model may be overfitting to specific tokens commonly found in classic problem statements. Since we observe many cases where state-of-the-art LLMs like GPT-4 Achiam et al. (2023) successfully identify logical fallacies under certain settings, we highlight the urgent need for a framework to tease out whether LLMs apply genuine reasoning or merely exploit token bias for their improved performance.
@@
Our study is unique from existing work Gou et al. (2023); Suri et al. (2024); Mukherjee and Chang (2024); Wang et al. (2024) in two folds. First, we are not evaluating the overall accuracy of LLMs in identifying different logical fallacies. Instead, our focus is on token bias. Although there are always more types of logical fallacies, we take the conjunction fallacy, syllogistic fallacy, and the "twenty-five horses" problem in graph theory as quintessential examples, which exhibit strong token biases that are more readily identifiable in their problem statements. By identifying and perturbing these specific tokens, we can induce predictable shifts in LLM responses. Second, we recognize that cognitive biases often emerge in implicit forms in real-life scenarios, so relying on engineering fine-grained prompts  Gou et al. (2023); Yao et al. (2024); Besta et al. (2024) to make LLMs identify specific logical fallacies is impractical for general-purpose user applications. As a result, we only leverage common prompting techniques that are sufficient to provide robust statistical evidence.
@@ 2 The General Framework
Once the underlying logic of a reasoning task is defined, we create an algorithm to generate a synthetic dataset with $n$ samples. While it is helpful to leverage LLMs for linguistic coherence in the process, the data generation should be carefully controlled, utilizing information from real-world data or established datasets to mitigate potential biases from purely AI-generated texts. The process begins with the creation of a curated list of entities, encompassing diverse names, genders, ages, occupations, cultural backgrounds, and events where applicable, along with a textual template that dictates the structure of the task description. By sampling from this list, we generate task descriptions that maintain the integrity and novelty of the dataset. This method ensures that while the LLM of interest might be familiar with the individual entities, it has never seen the specific combinations of these entities and narratives, thus bypassing data contamination.
@@ 3.1 Preliminaries
The syllogistic fallacy documents the logical failure that occurs when people are presented with syllogisms – two premises followed by a conclusion. Ideally, if the premises are true and the logical structure is valid, the conclusion must necessarily be true. However, when the argument’s structure is flawed, the conclusion may be invalid despite the surface-level logical form. Consider the following syllogism from Kahneman (2011): Is this logically sound?
@@ 3.2 Lost in Irrelevant Context
To further explain $\boldsymbol{H_{0}}$ and $\boldsymbol{H_{a}}$, the null hypothesis $\boldsymbol{H_{0}}$ always assumes that the LLM is a genuine reasoner and can consistently perform reasoning regardless of the superficial token changes, leading us to expect $\pi_{12}=\pi_{21}$. Meanwhile, token bias can systematically and predictably influence the LLM’s performance. Accordingly, $\pi_{12}<\pi_{21}$ aligns with our expectation that misleading tokens degrade performance, but an observation of $\pi_{12}>\pi_{21}$ where non-misleading options result in a greater decrease in performance - would be considered invalid.
@@ 3.3 Token Bias on Widely Cited Examples in Classic Literature
3.4 Token Bias on Well-Known Entity Names
@@
Token perturbation: Assume $P$ is a mathematical reasoning problem that involves a classic entity name in its story narratives. In contrast, the perturbed problem $P^{\prime}$ alters the classic name to another random one.
$\boldsymbol{H_{0}}$: $\pi_{12}=\pi_{21}$.
$\boldsymbol{H_{a}}$: $\pi_{12}>\pi_{21}$. ($\pi_{12}<\pi_{21}$ is invalid.)
@@
Here is another example of the token perturbation applied to the classic "twenty-five horses" problem in mathematical reasoning, as referenced in Figure 2. Note that perturbing the numbers is optional, but the total number of animals must always be a square number: There are $\rightarrow$ thirty-six $\rightarrow$ bunnies among which you need to find out the fastest three. You can conduct a race among at most $\rightarrow$ six to find out their relative speed. At no point, you can find out the actual speed of the $\rightarrow$ bunnies in a race. Find out the minimum number of races which are required to get the top $\rightarrow$ six $\rightarrow$ bunnies.
@@ 3.6 Leaking Hint Tokens
We also manually craft a detailed chain-of-thought instructions Wei et al. (2022b) that teach LLMs about correct reasoning steps and potential logical pitfalls, as shown in Appendix B.
@@ 4.2 Synthetic Dataset Generation
(c) Experimental results for Hypothesis with celebrity names 3.4 ($n=100$). The perturbed problems alternate the celebrity name to a generic one in problem statements. We run all different prompt methods. To reject the null, we expect ${\color[rgb]{0.921875,0.6796875,0.63671875}n12}<{\color[rgb]{0.28515625,0.5703125,0.6171875}n21}.$ We conclude that LLMs are frequently misled by celebrity names in problem statements that are irrelevant to logical essence.
@@ 4.3 Prompting Methods
We implemented commonly used prompting strategies that are sufficient for evaluating the null hypotheses within our framework. The specific prompting techniques we utilized are as follows, with their corresponding notations presented in Figure 4: Baseline: Directly answering the question without additional instructions. Zero-shot chain-of-thought (zs_cot): Includes the instruction "Let us think step by step" Wei et al. (2022b). One-shot (os): Involves a single in-context learning example Brown et al. (2020). Few-shots (fs): Utilizes three in-context examples. Similarly, we have os_cot and fs_cot. We also include weak_control_zs/os_cot and control_zs/os_cot for Hypothesis 3.6 representing prompts with additional weak or strong hints, as detailed in Appendix B.
@@ 5 Related Work
The Reasoning Capabilities of LLMs
There is an increasing number of works aim to improving the reasoning performance of LLMs Ouyang et al. (2022); Zhou et al. (2022); Lyu et al. (2023); Hao et al. (2023); Xu et al. (2024); Putta et al. (2024); Kumar et al. (2024); Yao et al. (2024); Cai et al. (2024), evaluating and critiquing their reasoning abilities Zhou et al. (2020); Hong et al. (2023); Huang et al. (2023); Shi et al. (2023); Sprague et al. (2024); Turpin et al. (2024); Xiao et al. (2024), trying to understand their reasoning processes Wei et al. (2022a); Merrill and Sabharwal (2023); Ma et al. (2023); Lanham et al. (2023); Merrill and Sabharwal (2024); Yang et al. (2024); Chen et al. (2024); Jin et al. (2024), as well as related surveys Qiao et al. (2022); Huang and Chang (2022); Ahn et al. (2024); Giadikiaroglou et al. (2024); Liang et al. (2024); Zheng et al. (2024) that comprehensively explore the reasoning capabilities of LLMs. Our work aligns with efforts that critique the generalization of LLMs’ reasoning capabilities, moving beyond accuracy-based benchmarks Talmor et al. (2018); Srivastava et al. (2022a, b); Fu et al. (2023); Yang et al. (2018), which primarily focus on overall question-answering accuracy, often without considering question augmentations or variations. Instead, we step one-level deeper into these problem statements, investigating potential token biases that may create the illusion of improved logical reasoning performance, while remaining vulnerable to irrelevant token perturbations. By examining these subtleties, our approach provides a more nuanced understanding of the limitations in LLM’s reasoning capabilities. Besides, we reformulate the evaluation as hypothesis testing, providing results with statistical guarantees.
@@
Recent studies Gardner et al. (2020); Hagendorff et al. (2023); Lin and Ng (2023); Talboy and Fuller (2023); Binz and Schulz (2023); Ullman (2023); Mitchell and Krakauer (2023) analyze the biases in LLMs with synthetic datasets. For example, Tamkin et al. (2023) uses an LLM to generate prompts that reveals patterns of discriminations in LLMs. Echterhoff et al. (2024) proposes a set of LLM-simulated experiments in a specific context. Although existing works Mukherjee and Chang (2024); Macmillan-Scott and Musolesi (2024); Wang et al. (2024); Suri et al. (2024); Jin et al. (2022) study more kinds of fallacy types in human psychology, they approach problems at a coarse level and only emphasize accuracy. Our study goes into a more fine-grained level with a series of hypotheses. We provide statistical guarantees and quantitative analyses of token bias that can be carefully and systematically tuned. Besides, Gou et al. (2023) presents the Rationality of Thought (RoT), decomposing responses into six predefined steps with hand-crafted prompt engineerings. Our work focuses on general prompting strategies that are sufficient to validate or reject our hypotheses.
@@ 6 Discussion
This work reconceptualizes the evaluation of the reasoning behavior of LLMs through the lens of token bias. The statistical evidence presented in our hypothesis-testing framework contributes to the larger discussion that LLMs do not always apply reasoning consistently in their decision-making processes. Instead, they primarily rely on token bias for response generation. This suggests that chain-of-thought prompting Wei et al. (2022b); Wang et al. (2022) or in-context learning Brown et al. (2020); Min et al. (2022); Lyu et al. (2022); Wang et al. (2022) may not elicit actual reasoning but instead result in semantic shortcuts for LLMs to imitate the desired behavior at superficial levels. These findings raise concerns about the extent to which LLMs truly engage in reasoning. Further investigations are needed to uncover the underlying mechanisms and limitations of LLMs’ reasoning capabilities.
@@ 7 Limitations
This hypothesis-testing framework is specifically designed for multiple-choice or yes/no questions and is not applicable to open-ended responses. It relies on LLMs with strong instruction-following capabilities to consistently produce responses that include the selected options, though we find that LLMs can generally follow these instructions in most cases. In addition, smaller LLMs, such as llama-3-8b-instruct, with lower instruction following capabilities may contain more confounders besides the token bias, which could weaken our hypothesis testing results. As a result, we mainly focus on state-of-the-art LLMs. Moreover, the finding of token biases requires manual efforts. We also acknowledge that there are likely other hypotheses and assumptions that a genuine reasoner should satisfy. Our current study focuses on the conjunction fallacy, syllogistic fallacy, the "twenty-five horses" problem in graph theory, and their variants to demonstrate our framework. These are quintessential examples and the framework could include a broader range of hypotheses, fallacy types, data modalities, and reasoning tasks.
8 Acknowledgement
This work was supported by NSF grant CCF-2112665 (TILOS), which provides funding for B.J. and C.J. It was also funded in part by the Laboratory Directed Research and Development (LDRD) program at Argonne National Laboratory, with support from the Office of Science, U.S. Department of Energy, under Contract No. DEAC02-06CH11357. Y. X. and W. J. Su acknowledge support from the NSF HDR TRIPODS award (CCF-1934876). The authors thank John K. Hutchison for his valuable suggestions regarding the motivation of this study.
@@ C.1 Conjunction Fallacy
Suppose Barack Obama is running for president in 2024. Which is more likely:
(a) Barack Obama will win the national popular vote
(b) Barack Obama will win the national popular vote but lose the Electoral College vote
@@ D.1 Hypothesis 1
The full experimental results for Hypothesis 3.2 are shown in Figure 5, 6 and Table LABEL:table:h1.
Table 2: Full Experimental results for Hypothesis 3.2. Note that in our experiments, $n^{*}=n_{21}+n_{12}$ is not equal to the number of data samples $n$. Here, $n_{12}$ denotes the instances where the LLM correctly answers the original problem but fails on the perturbed version, and $n_{21}$ denotes the opposite scenario. Thus, a large value of $n^{*}$ happens only when the LLM makes many mistakes. Specifically, GPT-4o in this table shows $n^{*}=1$ with few-shots learning. We find that GPT-4o is excellent in answering these problems with few-show exemplars, achieving near-perfect accuracy of almost $100\%$, so their $n_{12}$ and $n_{21}$ values are pretty small. This high accuracy, however, only lead us to fail to reject this particular instantiation of the null hypothesis, but not in other situations. While we could increase the sample size from the current $n$ to potentialy observe more errors, thus a higher $n^{*}$ in scenarios involving state-of-the-art LMs with few-shot learning, our rejection of the null hypothesis under other scenarios when tested against the GPT-4o is sufficient to argue that LLMs are not yet genuine reasoners.
@@
Figure 6: Comparison of the accuracy scores between the original and perturbed problems for Hypothesis 3.2.